OpenStudy (anonymous):

the objective is to solve systems of equations by the addition method (no multiplying required) problem: x + y = 5 3x - y = 7

OpenStudy (anonymous):

Just add the two equations and you get 4x = 12, so x = 3. Now plug this in, into either of the equations and you get y = 2.

OpenStudy (anonymous):

How would I show my work to this problem?

OpenStudy (anonymous):

x + y = 5 3x - y = 7 --------- 4x = 12 x = 12/4 = 3 substituting x = 3 in x + y = 5, you get 3 + y = 5, or y = 5 - 3 = 2.

OpenStudy (anonymous):

\[ x + y = 5 \ (1)\] \[ 3x - y = 7 \ (2)\] \[ (1)\ +\ (2) \implies 4x = 12 \implies x = 3\] I always solve simultanz by labelling the equations, but I'm cultured.